OpenStudy (anonymous):

HELP? I'll give a medal:) What type of line is the graph of 4x + y = -3? vertical horizontal rising falling

OpenStudy (jdoe0001):

let's graph it, can you solve it for "y"?

OpenStudy (anonymous):

y=-4x-3

OpenStudy (jdoe0001):

ok, let's get a couple of values x = 0 y = -4(0)-3 => y = -3 so if x =0, y = -3 (0, -3) x = 2 y = -4(2) -3 => y = -11 so if x = 2, y = -11 (2, -11)

OpenStudy (jdoe0001):

 when the graph from LEFT to RIGHT is going down, is so-called "falling" when the graph from LEFT to RIGHT is going up, is so-called "rising"

OpenStudy (anonymous):

so rising??/

OpenStudy (jdoe0001):

well, as you go from left to right, is "y" going UP or DOWN?

OpenStudy (anonymous):

down?

OpenStudy (jdoe0001):

down, right, so is "falling", the usual word for it is "decreasing"

OpenStudy (anonymous):

yes